40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sat Jan 21 01:20:28 2012, Frank, DailyProgress, RefCav, refcav work 
    Reply  Mon Jan 23 09:16:43 2012, rana, DailyProgress, RefCav, refcav work 
       Reply  Mon Jan 23 11:36:18 2012, Frank, DailyProgress, RefCav, refcav work 
Message ID: 791     Entry time: Sat Jan 21 01:20:28 2012     Reply to this: 793
Author: Frank 
Type: DailyProgress 
Category: RefCav 
Subject: refcav work 

re-did the RF wiring for both cavities:

  • Power splitters are now located on the table below the TTFSS box. This keeps all cables from/to the PDs and EOMs very short
  • Power level at res. EOM is ~20dBm, LO is ~7dBm (TTFSS has an internal amplifier) - Master LO is still a SR345 with 23dBm.
  • Removed all old cables which we don't need anymore and moved the ones used from the table to the framing structure to get less seismic coupling to the floating table from them.
  • Finished phase adjustment for the TTFSS and locked the cavity using all three actuators.
  • Temp feedback to the laser is now done by the TTFSS unit. There is NO computer control for any of the FSS stuff anymore - only for the PMC. cable to laser unit is temporary as we don't have the right cable length available.
  • Power is adjusted to 1mW (total)
  • We still don't have enough gain. Common gain is at max, fast at 782, (offset 500).
  • Measured in-loop ep noise floor - will add plot later. Noise floor measured at OUT1 of TTFSS box (common path, mixer output) is ~10nV/rtHz
  • Error signals are still pretty tiny (230 mVpp)
  • re-wired beat setup. Need to add a coupler to hook up the frequency counter in parallel. Will use that for characterization of radiative heater and temp feedback to it for beat note stabilization.

Next steps:

  • Full characterization of 14.75MHz res. PD from TNI
  • build a second PD for 14.75MHz.
  • add some channels to DAQ - there are currently no FSS channels at all
  • measure beat note by Tue
  • finish in-vac wiring
  • open vacuum can and install rad shields+heater
  • while in air measure TF from table to top stack with small accelerometer and compare with model
ELOG V3.1.3-